OpenStudy (anonymous):

Factor completely: 2x2 − 4x − 48

OpenStudy (anonymous):

@whpalmer4?

OpenStudy (whpalmer4):

\[2x^2-4x-48\] So what do you think we should do first? Do you see any common factors in all of those terms?

OpenStudy (anonymous):

divide all of the terms by 2

OpenStudy (whpalmer4):

Okay, that's a good start. \[2(x^2-2x-24)\]Right?

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

is it 2(x − 6)(x + 4)?

OpenStudy (whpalmer4):

yesh? have you been imbibing? :-)

OpenStudy (anonymous):

no I just couldint figure the first part out! Thanks!

OpenStudy (whpalmer4):

well, let's see! \[(x-6)(x+4) = x(x+4) - 6(x+4) = x^2 + 4x -6x -24 = x^2 -2x -24\] so if we multiply that by 2, we get our starting equation, good job!

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

@Loujoelou did you need something?

OpenStudy (whpalmer4):

just watching the exciting sport of competitive factoring :-)

OpenStudy (whpalmer4):

coming soon to a television channel near you!

OpenStudy (whpalmer4):

at about 4:30 AM, I think :-)

OpenStudy (anonymous):

XD :P it is very competitive!

OpenStudy (anonymous):

ha ha lol!

OpenStudy (anonymous):

Lol no thx tho, was just watching :p

OpenStudy (anonymous):

so whpalmer was right!